OpenStudy (anonymous):

4-8(x-11) -12=----------- -3 whats x and howd you find it?

OpenStudy (anonymous):

x=7 multiply both sides by -3 to get 36=4-8(x-11) 32=-8(x-11) x-11=-4 x=7